OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line (8,9); x+3y=7

OpenStudy (anonymous):

x+3y=7 3y=-x+7 y=-x/3 +7/3 (8,9), 9=-8/3 +c c=9+8/3 c=35/3 the req. equation is y=-x/3 +35/3 ???